OpenStudy (anonymous):

hey everyone i have a question related to advance geometric series , mind giving it a try plz :)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

attach the file by clicking "attach file" under the text box

OpenStudy (anonymous):

thnx :)

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

i know the partial sum formula is used

jimthompson5910 (jim_thompson5910):

so you'll use the formula \[\Large \sum_{i=1}^{n} a*r^{n-1} = a*\frac{1-r^n}{1-r}\]

OpenStudy (anonymous):

yes but i am finding difficult to write it down properly how ssould i obtain e

jimthompson5910 (jim_thompson5910):

sorry I should have used i instead of n in the exponent \[\Large \sum_{i=1}^{n} a*r^{\color{red}{\Huge i}-1} = a*\frac{1-r^n}{1-r}\]

jimthompson5910 (jim_thompson5910):

let me think

OpenStudy (anonymous):

so how would i get ratio , i think its e^(1/n)

OpenStudy (anonymous):

sure :)

jimthompson5910 (jim_thompson5910):

let's write out the few terms of this summation \[\LARGE \sum_{i=1}^{n} e^{i/n} = e^{1/n}+e^{2/n}+e^{3/n}+\ldots e^{n/n}\] what's the first term? what's the common ratio?

OpenStudy (anonymous):

the first term is e^1/n and the common ratio also e^1/n

jimthompson5910 (jim_thompson5910):

`the first term is e^1/n` correct

jimthompson5910 (jim_thompson5910):

`common ratio also e^1/n` incorrect

jimthompson5910 (jim_thompson5910):

pick any term you want divide it by the previous term

jimthompson5910 (jim_thompson5910):

so for example \[\Large r = \frac{e^{2/n}}{e^{1/n}} = ??\]

OpenStudy (anonymous):

e^1/n

OpenStudy (anonymous):

if u minus 2/n -1/n you get 1/n

jimthompson5910 (jim_thompson5910):

oh sorry I was thinking of something else, yes you are correct

jimthompson5910 (jim_thompson5910):

so a = e^(1/n) is the first term r = e^(1/n) is the common ratio they coincidentally are the same expression, but this isn't always the case

OpenStudy (anonymous):

lucky me :)

jimthompson5910 (jim_thompson5910):

now plug that into \[\Large a*\frac{1-r^n}{1-r}\]

OpenStudy (anonymous):

ok \[\frac{ e ^{1/n} (1-e) }{ (1-e ^{1/n})}\]

OpenStudy (anonymous):

thats it

jimthompson5910 (jim_thompson5910):

exactly \[\Large a*\frac{1-r^n}{1-r}\] \[\Large e^{1/n}*\frac{1-\left(e^{1/n}\right)^n}{1-e^{1/n}}\] \[\Large e^{1/n}*\frac{1-e^{(1/n)*(n/1)}}{1-e^{1/n}}\] \[\Large e^{1/n}*\frac{1-e^{n/n}}{1-e^{1/n}}\] \[\Large e^{1/n}*\frac{1-e^{1}}{1-e^{1/n}}\] \[\Large \frac{e^{1/n}(1-e)}{1-e^{1/n}}\]

OpenStudy (anonymous):

thank you so much :)

jimthompson5910 (jim_thompson5910):

no problem